OpenStudy (anonymous):

Find the coordinates of the midpoint of the segment whose endpoints are H(8, 2) and K(6, 10).

OpenStudy (anonymous):

7,6

OpenStudy (anonymous):

Take the averages of the x's and the y's i.e. what is the average x coordinate and what is the average y coordinate?

OpenStudy (anonymous):

@akash809 http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

mid point x=(x1+x2)/2 and y=(y1+y2)/2

OpenStudy (anonymous):

u a mod

OpenStudy (anonymous):

posted the concept as well

OpenStudy (anonymous):

No, but they thank me when I do that.

OpenStudy (anonymous):

I posted the concept. You posted the formula.

OpenStudy (anonymous):

@Kassandrac did u get it

OpenStudy (anonymous):

Yes, thank you so much.

OpenStudy (anonymous):

u r welcome